Baddie156:

What is an equation of the line that passes through the point (-6, – 7) and is perpendicular to the line 6x + 5y = 30?

daddysbabygirl:

6x + 5y = 30 5y = 30 - 6x y = 6 - (6/5) x So slope = -6/5 Perpendicular line has slope 5/6 (product of the slopes is -1) So y = 5/6 x + b Passed through (-6, -7) so -7 = -5 + b Sp b = -2 So y = (5/6)x - 2

Baddie156:

thx y'all 😘😘😘

daddysbabygirl:

np beautiful

Baddie156:

lol